OpenStudy (anonymous):

What is the minimum value of the pullup resistor RPuI (in Ohms) for which this inverter can obey the required static discipline?

OpenStudy (anonymous):

Vout=Vs*RON/(RON+RPULLUP). Vout is lower than VOL rpull= 4*r0n for inverter , nor rpull= 4*r0n for nand